# ECAMP: Entity-centered Context-aware Medical Vision Language Pre-training

### 20 Dec 2023 v1

@@ Abstract
[…]
@@ 1 Introduction
Cutting-edge algorithms [35] in medical vision-language pre-training (Med-VLP) primarily leverage the information-rich medical report-image pairs to learn generic representations from two aspects: contrastive-based methods align the text and image features at both global [13] and local [7, 12] levels; while reconstruction-based methods generate the masked words [2] and image patches [24] using cross-modality information [3]. Nevertheless, these methods still exhibit limitations in harnessing the potential of supervision in the report context. First, the linguistic challenge [38] posed by complex biomedical context [39] impeds the effectiveness of pre-training [1]. Second, radiologists often include diagnoses for common diseases (e.g., pneumothorax) in their reports, even when these diseases are negative [40]. This results in a significant context imbalance issue [1]: the majority of disease diagnoses are negative [^1] Statistical results of the imbalanced diagnosis descriptors can be found in the supplementary materials.”, introducing an adverse bias into the optimization of masked language modeling (MLM) [2].
@@ 2 Related work
2.1 General Vision-language Pre-training
@@ 2.2 Medical Vision Language Pre-training
Recent Med-VLP methods can be categorized into two distinct types: report-supervised cross-modal alignment pre-training and reconstruction-based self-supervised pre-training. The groundbreaking works of the former type such as ConVIRT [13], REFERS [27], and CheXZero [26] pre-train the model by directly maximizing mutual information between the global representations. GLoRIA [7], SAT [37], and MGCA [12] propose to align fine-grained features of paired image patches and words. Med-Unic [34], PRIOR [36], and BioVIL [1] attempt to comprehend intricate medical reports. MedKLIP [28] and KAD [29] extract the medical-related information using a triplet extraction module as additional supervised signals. Inspired by BERT [2] and MAE [24], the latter type learns representation by cross-modal context reconstruction task. M3AE [30] utilizes cross-attention for integrating multi-modal information to reconstruct masked tokens. While MRM [3] directly applies global average pooling for fusing visual features for masked text modeling. Further, MPMA [31] introduces a memory-augmented cross-modal fusion module to fully integrate visual information to assist report reconstruction.
Advanced efforts try to combine contrastive-based and reconstruction-based pertaining. CMITM [49] contrastively aligns global report and image representations after pre-training with MRM [3], while MedIM [33] boosts MGCA [12] with attention guided masked image modeling. Notwithstanding the enhancements, the aforementioned approaches are constrained by the intricacies of comprehending biomedical reports and the imbalance in diagnoses, which has spurred our pursuit of developing ECAMP.
@@ 2.3 Large Language Models (LLMs)
Method ChestX-ray14 (AUC) CheXpert (AUC) RSNA (AUC) COVIDx (ACC)
1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100%
Random init 60.0 65.2 72.8 70.4 81.1 85.8 71.9 82.2 88.5 64.2 75.4 87.7
CNN-based
GLoRIA [7] 70.7 78.6 84.1 86.6 87.8 88.1 86.1 88.0 88.6 73.3 86.8 90.5
${}^{{\dagger}}$ [37] 67.4 79.3 83.7 86.3 88.0 88.2 86.7 87.5 89.0 75.0 88.5 94.3
PRIOR [36] 75.7 79.4 84.3 86.2 88.3 88.6 85.7 87.1 89.2 79.5 89.8 94.8
MedKLIP# [28] 77.2 78.9 83.2 - - - 87.3 88.0 89.3 78.3 88.8 95.3
KAD [29] 78.7 80.7 82.5 - - - 89.8 91.8 92.5 78.5 90.5 95.8
ViT-based
MAE [24] 74.7 81.3 85.1 80.7 86.0 86.7 84.2 89.6 91.3 69.8 82.3 90.8
GLoRIA [7] 77.7 82.8 85.0 86.5 87.5 87.8 89.7 91.2 92.1 76.8 91.8 94.8
REFERS [27] 76.7 80.9 84.7 87.2 88.1 88.2 89.4 91.6 92.7 - - -
MGCA# [12] 78.2 82.7 85.0 87.0 88.4 88.5 90.7 92.6 93.4 75.2 91.5 94.3
MRM [3] 79.4 84.0 85.9 88.5 88.5 88.7 91.3 92.7 93.3 78.0 90.3 92.5
ECAMP (Ours) 81.2 84.5 86.5 88.7 88.9 89.2 91.5 92.9 93.8 81.0 93.5 96.3
[…]
@@ 3 Method
In this section, we first introduce how ECAMP distills distinct context from LLM in Sec. 3.1. Then, we illustrate the basic masked image modeling (MIM) [24] and the proposed entity-centered context-enhanced MLM in Sec. 3.2. Next, the process of context-guided super-resolution is shown in Sec. 3.3. Finally, we clarify the detailed approaches of multi-scale context fusion in Sec. 3.4 to integrate both global and local vision features for MLM.
@@ 3.1 Context Distillation
Given an image $I^{o}\in\mathbb{R}^{H^{o}\times W^{o}}$ and its original paired report $T_{o}$, Our target is to extract concise diagnosis for diseases from $T_{o}$. Accordingly, we select the set of entities $R=\{r^{1},r^{2},\ldots,r^{k}\}$ from MeSH [43], representing the diseases of interest in chest radiography. For each original report $T_{o}$, we identify entities $\mathbf{r}_{o}$ mentioned in the report through string matching. Next, prompt $p$ is designed to require LLM to comprehend the complex biomedical context and summarize the diagnosis for disease entities $\mathbf{r}_{o}$. Furthermore, to make the summarization of LLM stable, we leverage in-context learning [48] to construct a one-shot [47] sample $p^{\prime}$ in the final prompt [^2] The selected entities of disease and detailed prompt can be found in the Appendix.B and Appendix.C, respectively. $p=P(T_{o},\mathbf{r}_{o},p^{\prime})$. The distilled report $T_{d}$ is finally computed as follows:
@@ 3.2 Masked Image and Language Modeling
According to the statistical results in Appendix.A, the most negative descriptor “there is no” occurs approximately 20×20$\times$20 × more often than the most positive descriptor “there may be.”, we set $\lambda_{neg}=0.05$. Please note that $\lambda_{oth}$ is important to make sure that the model pays an enough attention to accurately reconstruct the critical “descriptive words” $T_{a}$. The final loss of entity-centered context-aware MLM is defined:
@@ 3.3 Context-guided Super-resolution
$\mathcal{L}=\mathcal{L}_{\mathrm{MIM}}+\mathcal{L}_{\mathrm{MLM}}+\mathcal{L}_{\mathrm{SR}},$ (6)
@@ 3.4 Multi-scale Context Fusion
Different from MRM [3] which directly adds the vision feature $f_{v}$ to the text embedding $E_{t}$, we first apply a self-attention (SA) for an initial comprehension of text embedding $f_{t}=\mathrm{SA}(E_{t})$. In order to simultaneously attain representative global and local features for multi-scale downstream tasks, we leverage patch-wise vision features $f_{v}$ as local vision features $f_{v}^{l}$ and use global average pooling (GAP) to compute global vision feature $f_{v}^{g}=\mathrm{GAP}(f_{v}^{l})$. Next, we aggregate local vision context $f_{v}^{l}$ with text feature $f_{t}$ by cross-attention (CA) to generate local aggregated feature $f_{a}^{l}=\mathrm{CA}(f_{v}^{l},f_{t})$, and project the $f_{v}^{g}$ by a linear projector to obtain global feature $f_{a}^{g}$. Finally, we integrate text feature $f_{t}$, local aggregated feature $f_{a}$, and duplicate global feature $f_{a}^{g}$ by addition.
@@ 4.2 Downstream Tasks Datasets
ChestX-ray14 [23] comprises 112,120 frontal-view chest radiographs, and introduces a multi-label classification task involving 14 common chest pathologies. The dataset is officially split into training, validation, and testing subsets with a ratio of 70%/10%/20%.
CheXpert [21] includes 223,648 chest radiographs from a frontal or lateral view, which categorizes each image into 5 diseases: atelectasis, cardiomegaly, consolidation, edema, and pleural effusion. Aligned with [13, 7], we designate the expert-labeled official validation set as the test data. We randomly chose 5,000 radiographs from the training data as the validation set. The split for training/validation/testing comprises 218,414/5,000/234 images.
RSNA Pneumonia [18] is composed of 29,700 frontal-view chest radiographs, each accompanied by pneumonia opacity bounding boxes if pneumonia is present in the image. Four tasks are conducted on this dataset: fine-tune classification, linear classification, semantic segmentation, and object detection. The official data split is followed, with the training/validation/test set consisting of 25,184/1,500/3,000 images, respectively.
@@ 4.4 Implementation Details
We utilize an 8-layer ViT-B/16 [11] as the backbone for the image encoder and a 4-layer ViT-B/16 for the image decoder. Additionally, we employ BERT [2] as the text encoder. If not otherwise stated, we resize images to 224$\times$224 as input, except for super-resolution supervision we retain the 448$\times$448 resolution images. Texts are padded or truncated to a fixed length of 256 tokens. The latent dimension, denoted as D, is set to 768. The implementation is based on PyTorch [58]. For pre-training, our framework runs for 115 epochs on 4 NVIDIA A100 GPUs, taking approximately 20 hours. The training batch size is set to 256 for each GPU, and we utilize a gradient accumulation step of 8. Consequently, the effective batch size reaches 8192. We opt for AdamW [59] as our default optimizer, with a learning rate of 1.5e-4 and weight decay of 0.05. The values for $\beta_{1}$ and $\beta_{2}$ are set to 0.9 and 0.95, respectively. For the LLM in Eq. 1, we use ChatGPT API [41] to distill 10000 reports and fine-tune a Vicuna model [60] to distill the rest reports.
@@ 4.5 Baselines
We conduct a comprehensive comparative analysis of ECAMP with numerous SOTA Med-VLP approaches using ResNet-50 [10] (‘CNN-Based’) or ViT-B/16 [11] (’ViT-Based’) as backbone, including MAE [24] trained on ImageNet [20] (denoted as ‘ImageNet init’), GLoRIA [7], REFERS [27], SAT [37], PRIOR [36], MedKLIP [28], MGCA [12], KAD [29], MedIM [33], Med-UniC [34], and MRM [3]. In addition, we pre-train MAE [24] (denoted as ‘MAE’) and GLoRIA [7] with ViT-B/16 on MIMIC-CXR [16] using official source code. For MGCA, MRM, SAT, PRIOR, KAD, and MedKLIP whose official model are released, we fine-tune their models on the public downstream datasets in case there is no corresponding performance metrics in referenced papers. For MedIM and Med-UniC, whose official models are absent, we directly copy the performances from their papers for comparison.
@@ 5.1 Results on Classification
We conduct both fine-tune and linear-probe with our ECAMP on four datasets, using 1%, 10%, and 100% of the training data, respectively. As in Table 1, our ECAMP significantly outperforms all CNN and ViT-based competitive Med-VLP models across different ratios of training data on the four datasets. It is worth noting that, on the ChestX-ray14, which encompasses the most of testing data and disease labels, ECAMP surpasses the strongest counterpart by 1.8% using only 1% labeled data. For a comparison of 14 diseases, refer to Appendix.D. Surprisingly, on the COVIDx with a novel disease “COVID-19”, which is not present in MIMIC-CXR [16], a great performance gap can be found in Table 1 between ECAMP and SOTA methods. This validates the superiority of the representation learned by our ECAMP. Interestingly, we find the cross-modality reconstruction-based methods MRM [3] and our ECAMP outperform the contrastive-based methods like MGCA [12] and GloRIA [7]. Moreover, Table 2 presents the results of linear-probe classification. ECAMP exhibits consistently improved performance compared with SOTA methods on ChestX-ray 14 and COVIDx datasets. However, there is a limitation of ECAMP for linear probing on CheXpert.
Context distillation Descriptors masking Re-balanced factor Super resolution MSCF ChestX-ray14 (AUC) COVIDx (ACC) SIIM (Dice) RSNA (mAP)
Global Local
$\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ 79.5 78.0 65.4 4.1
$\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ 80.0 81.0 66.2 9.7
$\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ 80.2 81.5 66.4 13.6
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ 80.7 82.0 66.5 12.5
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 81.4 83.3 65.6 7.6
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ 80.6 81.5 66.9 19.0
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 81.2 84.3 66.7 17.5
Table 4: Ablation study of five simple-yet-effective components (context distillation, descriptors masking, re-balanced factor, context-guided super-resolution, and MSCF) proposed in ECAMP on multi-scale downstream tasks including fine-tune classification (ChestX-ray14 [23]), linear classification (COVIDx [22]), semantic segmentation (SIIM [19]), and object detection (RSNA [18]). All of the downstream tasks are trained with 1% training data.
$\lambda_{neg}$ ChestX-ray14 (AUC) COVIDx (ACC)
@@
0.025 80.9 84.3 80.0 89.3
0.05 81.2 84.5 84.3 91.8
0.1 80.9 84.2 80.5 89.5
0.25 80.5 83.9 80.8 91.3
0.5 80.7 84.1 82.5 85.8
Table 5: Ablation study on parameter $\lambda_{neg}$ proposed in Eq. 3 on fine-tuning (ChestX-ray14 [23]) and linear (COVIDx [22]) classification with different ratios of training data (1%, 10%).
5.2 Results on Detection and Segmentation
To assess the effectiveness of local representations learned by ECAMP, extensive experiments on fine-grained downstream tasks including segmentation and detection are conducted in Table. 3. Compared to strong SOTA approaches and boosted by MSCF, ECAMP makes remarkable improvements on all three datasets and ratios. Even when MGCA [12] and MedKLIP [28] leverage disease-level labels, and Med-UniC [34] employs additional large-scale training data like PadChest [17]. Distinguished Dice score leaps of 4.6% and 1.0% against SOTA methods are achieved on two segmentation datasets SIIM [19] and RSNA [18] using 1% training data, respectively. Notably, ECAMP reaches a Dice score of 66.7%, which outperforms CNN-based methods fine-tuned by 100% training data on SSIM. Similarly, our method attains the best mAP using limited data on the object detection task. These achievements substantiate the effectiveness of ECAMP to optimize informative local features for fine-grained downstream tasks.

### 18 Mar 2024 v2

@@ Abstract
Keywords:
medical vision-language pre-training masked modeling
@@ 1 Introduction
Cutting-edge algorithms [1] in medical vision-language pre-training (Med-VLP) primarily leverage the information-rich medical report-image pairs to learn generic representations from two aspects: contrastive-based methods align the text and image features at both global [61] and local [17, 49] levels; while reconstruction-based methods generate the masked words [11] and image patches [15] using cross-modality information [63]. Nevertheless, these methods still exhibit limitations in harnessing the potential of supervision in the report context. First, the linguistic challenge [22] posed by complex biomedical context [29] impeds the effectiveness of pre-training [2]. Second, radiologists often include diagnoses for common diseases (e.g., pneumothorax) in their reports, even when these diseases are negative [10]. This results in a significant context imbalance issue [2]: the majority of disease diagnoses are negative [^1] Statistical results of the imbalanced diagnosis descriptors can be found in the supplementary materials., introducing an adverse bias into the optimization of masked language modeling (MLM) [11].
@@ 2 Related Work
2.1 General Vision Language Pre-training
@@ 2.2 Medical Vision Language Pre-training
Recent Med-VLP methods can be categorized into two distinct types: report-supervised cross-modal alignment pre-training and reconstruction-based self-supervised pre-training. The groundbreaking works of the former type such as ConVIRT [61], REFERS [62], and CheXZero [47] pre-train the model by directly maximizing mutual information between the global representations. GLoRIA [17], SAT [31], and MGCA [49] propose to align fine-grained features of paired image patches and words. PRIOR [7] and BioVIL [2] attempt to comprehend intricate medical reports. MedKLIP [54] and KAD [60] extract the medical-related information using a triplet extraction module as additional supervised signals. Inspired by BERT [11] and MAE [15], the latter type learns representation by cross-modal context reconstruction task. M3AE [6] utilizes cross-attention for integrating multi-modal information to reconstruct masked tokens. While MRM [63] directly applies global average pooling for fusing visual features for masked text modeling. Further, MPMA [57] introduces a memory-augmented cross-modal fusion module to fully integrate visual information to assist report reconstruction.
Advanced efforts try to combine contrastive-based and reconstruction-based pre-training. CMITM [5] contrastively aligns global report and image representations after pre-training with MRM [63], while MedIM [56] boosts MGCA [49] with attention guided masked image modeling. Notwithstanding the enhancements, the aforementioned approaches are constrained by the intricacies of comprehending biomedical reports and the imbalance in diagnoses, which has spurred our pursuit of developing ECAMP. Med-Unic [48] and Med-MLLM [32] further improve the performance by leveraging more pre-training datasets.
@@ 3 Method
In this section, we first introduce how ECAMP leverages context distillation to distill distinct context from LLM in Sec. 3.1. Then, we illustrate the basic masked image modeling (MIM) [15] and the proposed entity-centered context-enhanced MLM in Sec. 3.2. Next, Sec. 3.3 shows the process of context-guided super-resolution. Finally, we clarify the detailed approaches of multi-scale context fusion in Sec. 3.4 to integrate both global and local vision features for MLM.
@@ 3.1 Context Distillation
Given an image $I^{o}\in\mathbb{R}^{H^{o}\times W^{o}}$ and its original paired report $T_{o}$, Our target is to extract concise diagnosis for diseases from $T_{o}$. Accordingly, we select the set of entities $R=\{r^{1},r^{2},\ldots,r^{k}\}$ from MeSH [30], representing the diseases of interest in chest radiography. For each original report $T_{o}$, we identify entities $\mathbf{r}_{o}$ mentioned in the report through string matching. Next, prompt $p$ is designed to require LLM to comprehend the complex biomedical context and summarize the diagnosis for disease entities $\mathbf{r}_{o}$. Furthermore, to make the summarization of LLM stable, we leverage in-context learning [34] to construct an one-shot [53] sample $p^{\prime}$ in the final prompt [^2] The selected entities of disease and detailed prompt can be found in the Appendix.B and Appendix.C, respectively. $p=P(T_{o},\mathbf{r}_{o},p^{\prime})$. The distilled report $T_{d}$ is finally computed as follows:
@@ 3.2 Masked Image and Language Modeling
According to the statistical results in Appendix.A, the most negative descriptor “there is no" occurs approximately 20×20$\times$20 × more often than the most positive descriptor “there may be.", we set $\lambda_{neg}=0.05$. Please note that $\lambda_{oth}$ is important to make sure that the model pays an enough attention to accurately reconstruct the critical “descriptive words” $T_{a}$. The final loss of entity-centered context-aware MLM is defined:
@@
[…]
Method CXR14 (AUC) CXP (AUC) RSNA (AUC) COVIDx (ACC)
1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100%
Random init 60.0 65.2 72.8 70.4 81.1 85.8 71.9 82.2 88.5 64.2 75.4 87.7
CNN-based
GLoRIA [17] 70.7 78.6 84.1 86.6 87.8 88.1 86.1 88.0 88.6 73.3 86.8 90.5
${}^{{\dagger}}$ [31] 67.4 79.3 83.7 86.3 88.0 88.2 86.7 87.5 89.0 75.0 88.5 94.3
PRIOR [7] 75.7 79.4 84.3 86.2 88.3 88.6 85.7 87.1 89.2 79.5 89.8 94.8
MedKLIP# [54] 77.2 78.9 83.2 87.1 88.4 88.6 87.3 88.0 89.3 78.3 88.8 95.3
KAD [60] 78.7 80.7 82.5 87.2 88.2 88.5 89.8 91.8 92.5 78.5 90.5 95.8
ViT-based
MAE [15] 74.7 81.3 85.1 80.7 86.0 86.7 84.2 89.6 91.3 69.8 82.3 90.8
GLoRIA [17] 77.7 82.8 85.0 86.5 87.5 87.8 89.7 91.2 92.1 76.8 91.8 94.8
REFERS [62] 76.7 80.9 84.7 87.2 88.1 88.2 89.4 91.6 92.7 76.5 90.8 94.5
MGCA# [49] 78.2 82.7 85.0 87.0 88.4 88.5 90.7 92.6 93.4 75.2 91.5 94.3
MRM [63] 79.4 84.0 85.9 88.5 88.5 88.7 91.3 92.7 93.3 78.0 90.3 92.5
ECAMP (Ours) 81.9 85.1 86.7 88.8 89.5 89.7 91.5 93.2 94.2 83.0 95.5 97.5
@@ 3.3 Context-guided Super-resolution
$\mathcal{L}=\mathcal{L}_{\mathrm{MIM}}+\mathcal{L}_{\mathrm{MLM}}+\mathcal{L}_{\mathrm{SR}}.$ (6)
@@ 3.4 Multi-scale Context Fusion
Different from MRM [63] which directly adds the vision feature $f_{v}$ to the text embedding $E_{t}$, we first apply a self-attention (SA) for an initial comprehension of text embedding $f_{t}=\mathrm{SA}(E_{t})$. In order to simultaneously attain representative global and local features for multi-scale downstream tasks, we leverage patch-wise vision features $f_{v}$ as local vision features $f_{v}^{l}$ and use global average pooling (GAP) to compute global vision feature $f_{v}^{g}=\mathrm{GAP}(f_{v}^{l})$. Next, we aggregate local vision context $f_{v}^{l}$ with text feature $f_{t}$ by cross-attention (CA) to generate local aggregated feature $f_{a}^{l}=\mathrm{CA}(f_{v}^{l},f_{t})$, and project the $f_{v}^{g}$ by a linear projector to obtain global feature $f_{a}^{g}$. Finally, we integrate text feature $f_{t}$, local aggregated feature $f_{a}^{l}$, and duplicate global feature $f_{a}^{g}$ by addition.
@@ 4.2 Downstream Tasks Datasets
ChestX-ray14 (CXR14) [51] comprises 112,120 frontal-view chest radiographs, and introduces a multi-label classification task involving 14 common chest pathologies. The dataset is officially split into training, validation, and testing subsets with a ratio of 70%/10%/20%.
CheXpert (CXP) [18] includes 223,648 chest radiographs from a frontal or lateral view, which categorizes each image into 5 diseases: atelectasis, cardiomegaly, consolidation, edema, and pleural effusion. Aligned with [61, 17], we designate the expert-labeled official validation set as the test data. We randomly chose 5,000 radiographs from the training data as the validation set. The split for training/validation/testing comprises 218,414/5,000/234 images.
RSNA Pneumonia [43] is composed of 29,700 frontal-view chest radiographs, each accompanied by pneumonia opacity bounding boxes if pneumonia is present in the image. Four tasks are conducted on this dataset: fine-tune classification, linear classification, semantic segmentation, and object detection. For fine-tune and linear classification tasks, we adhere to the official data partitioning, allocating 25,184, 1,500, and 3,000 images for the training, validation, and test sets, respectively. In the context of semantic segmentation and object detection, our approach aligns with the methodology proposed in [49], wherein the original training set is randomly divided into subsets of 16,010, 5,337, and 5,337 samples for training, validation, and testing purposes, respectively.
@@ 4.4 Implementation Details
We utilize an 8-layer ViT-B/16 [12] as the backbone for the image encoder and a 4-layer ViT-B/16 for the image decoder. Additionally, we employ BERT [11] as the text encoder. If not otherwise stated, we resize images to 224$\times$224 as input, except for super-resolution supervision we retain the 448$\times$448 resolution images. Texts are padded or truncated to a fixed length of 256 tokens. The latent dimension, denoted as D, is set to 768. The implementation is based on PyTorch [38]. For pre-training, our framework runs for 115 epochs on 4 NVIDIA A100 GPUs, taking approximately 20 hours. The training batch size is set to 256 for each GPU, and we utilize a gradient accumulation step of 8. Consequently, the effective batch size reaches 8,192. We opt for AdamW [33] as our default optimizer, with a learning rate of 1.5e-4 and weight decay of 0.05. The values for $\beta_{1}$ and $\beta_{2}$ are set to 0.9 and 0.95, respectively. The mean squared error (MSE) loss is employed for masked image modeling and super-resolution, while cross-entropy loss is used for masked language modeling.
For the LLM in Eq. 1, we use GPT-3.5-turbo API [36] with setting temperature to 0 to distill 10,000 reports and fine-tune a Vicuna model [8] to distill the rest reports.
@@ 4.5 Baselines
We conduct a comprehensive comparative analysis of ECAMP with numerous SOTA Med-VLP approaches pre-trained solely on MIMIC-CXR using ResNet-50 [16] (‘CNN-Based’) or ViT-B/16 [12] (‘ViT-Based’) as backbone, including MAE [15] trained on ImageNet [21] (denoted as ‘ImageNet init’), GLoRIA [17], REFERS [62], SAT [31], PRIOR [7], MedKLIP [54], MGCA [49], KAD [60], MedIM [56], and MRM [63]. Notably, Med-UniC [48] uses external multilingual datasets. In addition, we pre-train MAE [15] (denoted as ‘MAE’) and GLoRIA [17] with ViT-B/16 on MIMIC-CXR [19] using official source code. For MGCA, MRM, SAT, PRIOR, KAD, and MedKLIP whose official model are released, we fine-tune their models on the public downstream datasets in case there is no corresponding performance metrics in referenced papers. For MedIM and Med-UniC, whose official models are absent, we directly copy the performances from their papers for comparison.
Table 4: Ablation study of five simple-yet-effective components (context distillation (CD), descriptors masking (DM), re-balanced factor (RF), context-guided super-resolution (SR), and MSCF) proposed in ECAMP on multi-scale downstream tasks including fine-tune classification (CXR14 [51] ), linear classification (COVIDx [50]), semantic segmentation (SIIM [44]), and object detection (RSNA [43]). All of the downstream tasks are trained with 1% training data.
CD DM RF SR MSCF CXR14 (AUC) COVIDx (ACC) SIIM (Dice) RSNA (mAP)
Global Local
$\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ 79.5 78.0 65.4 4.1
$\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ 80.2 81.5 65.9 12.2
$\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ 81.1 82.3 66.5 16.1
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ 81.5 83.0 66.6 14.9
$\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 81.3 82.5 66.9 15.6
$\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 81.4 83.0 66.4 16.8
$\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 81.1 84.8 67.1 18.9
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 81.8 85.0 65.7 12.6
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ 81.4 83.3 67.5 20.4
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 81.9 86.3 67.3 19.8
@@ 5.1 Results on Classification
We conduct both fine-tune and linear-probe with our ECAMP on four datasets, using 1%, 10%, and 100% of the training data, respectively. As in Table 1, our ECAMP significantly outperforms all CNN and ViT-based competitive Med-VLP models across different ratios of training data on the four datasets.
Table 5: Ablation study on parameter $\lambda_{neg}$ proposed in Eq. 3 on fine-tuning (CXR14 [51]) and linear (COVIDx [50]) classification with different ratios of training data (1%, 10%).
$\lambda_{neg}$ CXR14 (AUC) COVIDx (ACC)
@@
0.025 81.7 84.5 85.0 93.3
0.05 81.9 85.1 86.3 95.0
0.1 81.8 84.6 80.5 94.8
0.25 81.5 84.2 79.5 94.0
0.5 81.4 84.2 82.5 91.5
It is worth noting that, on the ChestX-ray14 [51], which encompasses the most of testing data and disease labels, ECAMP surpasses the strongest counterpart by 2.5% using only 1% labeled data. For a comparison of 14 diseases, refer to Appendix.D. Surprisingly, on the COVIDx with a novel disease “COVID-19”, which is not present in MIMIC-CXR [19], a great performance gap can be found in Table 1 between ECAMP and SOTA methods. This validates the superiority of the representation learned by our ECAMP. Interestingly, we find the cross-modality reconstruction-based methods MRM [63] and our ECAMP outperform the contrastive-based methods like MGCA [49] and GloRIA [17]. Moreover, Table 2 presents the results of linear-probe classification. ECAMP exhibits consistently improved performance compared with SOTA methods on CXR14, RSNA and COVIDx datasets.
5.2 Results on Segmentation and Detection
To assess the effectiveness of local representations learned by ECAMP, extensive experiments on fine-grained downstream tasks including segmentation and detection are conducted in Table. 3. Compared to strong SOTA approaches and boosted by MSCF, ECAMP makes remarkable improvements on all three datasets and ratios. Even in scenarios where MGCA [49] and MedKLIP [54] leverage disease-level labels, and Med-UniC [48] incorporates additional large-scale training data like PadChest [4], our ECAMP consistently outperforms all aforementioned methods. Distinguished Dice score leaps of 4.2% and 1.5% against SOTA methods are achieved on two segmentation datasets SIIM [44] and RSNA [43] using 1% training data, respectively. Notably, ECAMP reaches a Dice score of 67.3%, which outperforms CNN-based methods fine-tuned by 100% training data on SIIM [44]. Similarly, our method attains the best mAP using limited data on the object detection task. Remarkably, our method exhibits a significant lead of 8.3% compared to other SOTA open-source methods based on ViT. These achievements substantiate the effectiveness of ECAMP to optimize informative local features for fine-grained downstream tasks.